A fluid at a pressure of 3 bar and specific volume of 0.18 m3/kgm contained in a piston cylinder arrangement expands reversibly to a pressure of 0.6 bar according to relation p = c/v2 where c is constant. Compute the work done in cal/kgm by fluid on the piston.


1
Expert's answer
2021-01-20T04:57:18-0500

pV2 = C

p1 = 3 bar = 3 x 105 N/m2

V1 = 0.18 m3/kg

W = area under pV graph

W = "\smallint" pdV = "\smallint" C/V2 dV = -C(1/V2 - 1/V1)

W = C(1/V1 - 1/V2)

C = pV2 = p1V12 = 3 x 0.182 = 0.0972 bar (m3/kg2)

V2 = (C/p2)1/2 = (0.0972/0.6)1/2 = 0.402 m3/kg

W = 0.0972 x (1/0.18 - 1/0.402) = 29840 Nm/kg = 7127 cal/kg